### v2

@@ 5 Conclusion
In this study, we trained a transformer model using either the MLM or the MTR objectives to predict six ADME endpoints (solubility, permeability, microsomal stability in both humans and rats, and plasma protein binding in both humans and rats). We experimented with different setups of pre-training dataset sizes, starting from a randomly initialized model to a model trained on the full GucaMol dataset of 1.3M molecules. We proposed domain adaptation as an extra training step where the pre-trained model is further trained on the molecules from the downstream task itself without providing labels. We also investigated the MTR and CL objectives to understand how chemically-aware objectives perform compared to a lexical objective like MLM. We showed that training a transformer model can be performed very efficiently when taking into account the data diversity and the appropriate training objectives. The most efficient, yet performant, model setup was pre-training with $\sim 400$K molecules using the MLM objective and domain adaptation using the endpoint molecules ($<3$K) using the MTR objective. Our model competes against more complicated models like MolBERT [1] (pre-trained on 1.3M molecules using MLM, MTR, and SMILES-Eq) and MolFormer [2] (pre-trained on 100M molecules using the MLM objective).
In conclusion, we highlighted the potential for developing efficient and high-performing molecular transformer models. We also highlighted areas of improvement like investigating pre-training data quality or diversity, identifying which objectives better suited for capturing the nuances of a property endpoint, and interrogating the architectural choices of the model.
@@ 6 Data and code availability
All data, analysis and implementation code can be found in our github repo at https://github.com/uds-lsv/domain-adaptation-molecular-transformers
@@
Figure S1: Data summary and distributions of the six investigated endpoint after removing censored labels
Figure S2: Significance analysis of the 0% vs 30% setups only. DataSAIL splitting is used
Figure S3: Significance analysis of the increased pre-training dataset size using DataSAIL splitting
Figure S4: Significance analysis of the 0% vs 30% setups only. Random splitting is used
Figure S5: The effect of increasing pre-training dataset size using random splitting.
Figure S6: Significance analysis of the increased pre-training dataset size using random splitting
Figure S7: The effect of domain adaptation with different sizes of pre-training dataset using random splitting.
Figure S8: Significance analysis of domain adaptation with different sizes of pre-training dataset using random splitting.
Figure S9: Significance analysis of domain adaptation across different sizes of pre-training dataset under the MTR objective. random splitting was performed.
Figure S10: Performance of the most efficient model from our analysis to models from the literature. 30% + MTR corresponds to a transformer model pre-trained with 30% of the GuacaMol dataset then domain adapted on the corresponding endpoint using the MTR objective. MolBERT [1] and MolFormer [2] are two transformer models from the literature. RF + PhysChem corresponds to a model trained on the physicochemical properties calculated by the RDKit package. The results are shown for the random splitting scheme.
Figure S11: Significance analysis of the literature comparison.
Figure S12: Physicochemical properties distributions and data clusters of the Solubility endpoint
Figure S13: Physicochemical properties distributions and data clusters of the permeability endpoint
Figure S14: Physicochemical properties distributions and data clusters of the HLM endpoint
Figure S15: Physicochemical properties distributions and data clusters of the RLM endpoint
Figure S16: Physicochemical properties distributions and data clusters of the hPPB endpoint
Figure S17: Physicochemical properties distributions and data clusters of the rPPB endpoint
Figure S18: Performance comparison between the models pre-trained with MLM vs MTR objectives. Random splitting was used.
Figure S19: The effect of increasing pre-training dataset size for the MTR-pre-trained models using DataSAIL splitting.
Figure S20: The effect of increasing pre-training dataset size for the MTR-pre-trained models using Random splitting.
Figure S21: Significance analysis of the increased pre-training dataset size with the MTR objectives using DataSAIL splitting
Figure S22: Significance analysis of the increased pre-training dataset size with the MTR objectives using random splitting
Figure S23: comparing pre-training and domain adaptation objectives to identify the best approach. the models names follow the scheme of pre-training objective_pre-training dataset size_domain adaptation objective. For example, MLM_30%_No DA corresponds to a model pre-trained with the MLM objective on 30% of the pre-training dataset and without domain adaptation. Random splitting was used
Figure S24: Significance analysis of the different models from figure S23. Random splitting was used
Figure S25: Significance analysis of the different models from figure 7. DataSAIL splitting was used
Figure S26: Comparison of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors. DataSAIL splitting is used.
Figure S27: Significance analysis of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors (figure S26.
Figure S28: Comparison of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors. random splitting is used.
Figure S29: Significance analysis of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors (figure S28).

### v3

@@ 5 Conclusion
In this study, we systematically investigated the performance of transformer-based models for seven molecular property datasets, evaluating the effects of pre-training strategies and domain adaptation objectives. Our results show that while large-scale pre-training with generic objectives like masked language modeling (MLM) offers some benefit, performance plateaus beyond a certain scale. In contrast, domain adaptation using a chemically informed multi-task regression (MTR) objective on domain molecules led to consistent and statistically significant improvements across diverse ADME datasets, even when applied to $\leq$ 4K molecules.
The most effective configuration was a model pre-trained on $\sim 400$K molecules (30% of the GuacaMol dataset) and hybrid objectives that combined MLM for pre-training and MTR for domain adaptation. This model (MLM_MTR) demonstrated competitive or superior performance to larger transformer models such as MolBERT [20] and MolFormer [26], despite being trained on significantly fewer molecules and lightweight objectives. Furthermore, our comparison with traditional machine learning baselines revealed that models explicitly leveraging physicochemical descriptors—like Random Forests or MTR-adapted transformers—outperformed purely structure-based approaches.
While the baseline Random Forest model using raw physicochemical properties remained the strongest overall, our findings highlight clear and practical strategies for enhancing the performance of transformer models. Specifically, incorporating chemically informed objectives, aligning model adaptation with task-relevant data, and using domain adaptation in addition to pre-training were all critical factors.
@@ 6 Data and code availability
All data, analysis and implementation code can be found in our github repository at https://github.com/uds-lsv/domain-adaptation-molecular-transformers. The models pre-trained with 30% and 60% using MLM or MTR can be found on our HuggingFace collection at https://huggingface.co/collections/UdS-LSV/domain-adaptation-molecular-transformers-6821e7189ada6b7d0a5b62d4. We additionally uploaded the domain-adapted models using MTR for each examined endpoint.
@@
Figure S1: Fang et. al. [3] dataset summary and distribution.
Figure S2: AstraZeneca ChEMBL [31] dataset summary and distribution.
Figure S3: MAE performance for increasing pre-training dataset size using random splitting. 0% corresponds to a randomly initialized model with no pre-training and 100% correspond to the $\sim$1.3M molecules of the GuacaMol dataset. Two-tailed significance analysis were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.
Figure S4: $R^{2}$ performance for increasing pre-training dataset size using random splitting. 0% corresponds to a randomly initialized model with no pre-training and 100% correspond to the $\sim$1.3M molecules of the GuacaMol dataset. Two-tailed significance analysis were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.
Figure S5: MAE performance for a baseline model trained with pre-training only (No DA), and three models incorporating domain adaptation (DA) using different objectives: Masked Language Modeling (MLM), Contrastive Learning (CL), and Multi-task Regression (MTR) for physicochemical properties. P-values are from one-tailed paired t-tests comparing each DA model to the No DA baseline, under the hypothesis that DA improves performance. Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$. Data is randomly split.
Figure S6: $R^{2}$ performance for a baseline model trained with pre-training only (No DA), and three models incorporating domain adaptation (DA) using different objectives: Masked Language Modeling (MLM), Contrastive Learning (CL), and Multi-task Regression (MTR) for physicochemical properties. P-values are from one-tailed paired t-tests comparing each DA model to the No DA baseline, under the hypothesis that DA improves performance. Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$. Data is randomly split.
Figure S7: $R^{2}$ performance of the most efficient model from our analysis to models from the literature. MLM_MTR corresponds to a transformer model pre-trained with $\sim 400$K molecules using the MLM objective then domain adapted on the corresponding endpoint using the MTR objective. MolBERT [20] is a chemically aware transformer than has been pre-trained on $\sim 1.3$M molecules using MLM, MTR, and SMILES-EQ objectives. MolFormer [26] is a large-scale transformer pre-trained on 100M molecules using MLM. PhysChem and Morgan fingerprint correspond to two baselines using Random Forest models. Two-tailed significance analysis were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.